# FocusLLM: Scaling LLM’s Context by Parallel Decoding

### v1

@@ 3 Experiments
In this section, we will conduct a comprehensive evaluation of the effectiveness of FocusLLM, spanning both language modeling and a variety of downstream tasks.
3.1 Experimental Details
We aligned most of our experimental settings with those of Activation Beacon Zhang et al. (2024a) to ensure comparable results. Specifically, we conducted training on a Linux server equipped with 8×A100 GPUs, each with 40GB of memory. The training was carried out for 10,000 steps, equivalent to one epoch of the entire training dataset, using a batch size of 8 and a learning rate of 5e-5 with a linear scheduler. To conserve GPU memory, we employed deepspeed’s zero2_offload optimizing stage. The training process was completed in approximately 20 hours.
For hyper-parameters, during training, the chunk size was randomly selected from the set {64, 128, 256, 1024, 2048}. For the length of tokens injected into each chunk, we set a default of 512 tokens for inference. And we ensured this length did not exceed the chunk size in the training procedure. As a result, the length of injected tokens was $\min\{512,chunk\ size\}$.
@@ 3.2 Long-context Language Modeling
In this section, we evaluate the fundamental capabilities of FocusLLM on long-context language modeling benchmarks, with text lengths ranging from 4K to 128K tokens.
@@
, and the training-free method StreamingLLM Xiao et al. (2023), which is based on attention sinks. ii) Fine-tuned methods trained on long inputs, such as LongAlpaca-16K Chen et al. (2023b), LongChat-32K Li et al. (2023a), and YaRN-128K Peng et al. (2023). iii) Methods with designed structures specifically for long contexts, including AutoCompressor-6K Chevalier et al. (2023), LongLlama Tworkowski et al. (2024) and Activation Beacon Zhang et al. (2024a). For instance, Activation Beacon achieves compression of long texts by training the model to represent the information of a regular text segment with a small number of beacon tokens.
Analysis. The results are presented in Table 2. Here are several observations we can make: (1) Compared to the basic LLaMA-2-7B model and some fine-tuning free methods, our model demonstrates superior performance. When extending the context length from 4K to longer, the perplexity becomes lower, indicating that information from a longer context can be effectively utilized. (2) There is a slight increase in the perplexity at longer lengths on CodeParrot. One possible reason is that as the number of candidate tokens from parallel decoding in the memory increases, the model’s output probabilities become more evenly distributed. However, this does not affect the effectiveness in downstream tasks (as discussed in Section 3.3). (3) FocusLLM achieves comparable performance to fine-tuned full-attention methods. This result is notable because our model operates with significantly higher training efficiency. For instance, LongLlama is fine-tuned using 7B tokens with all parameters being trainable. In contrast, FocusLLM uses $1/10$ of the training budget and $1/3$ of the parameters. (4) FocusLLM can maintain language modeling capabilities at lengths much longer than other models while retaining precise comprehension of the entire text. In experiments, FocusLLM can maintain stable perplexity at lengths of 400K, which means a 100-fold expansion of the original LLaMA-2-7B’s context length. Although models like StreamingLLM and Activation Beacon can still achieve lower perplexity by compressing tokens, they are unable to recover the previous context information, which severely affects their capabilities in downstream tasks.
In summary, FocusLLM achieves comparable language modeling performance on shorter sequences with a small training cost, while also being capable of handling much longer texts. Building on the design of parallel decoding, FocusLLM avoids the information loss associated with the preceding context, which is a common issue in compression models. We will provide a detailed demonstration of the downstream task performance in the subsequent sections.
Table 3: The results of Vicuna-based models on $\infty$-Bench and LongBench. The models on the right part can process extremely long inputs. On both benchmarks, FocusLLM achieves significant improvements compared to strong baselines.
@@
$\infty$-Bench (214k tokens) Math.Find 11.71 9.43 13.43 17.14 OOM OOM 6.00 11.14 11.71
@@ 3.3 Downstream Tasks
Datasets. To assess the capabilities of FocusLLM in real-world scenarios, we select two widely used datasets: Longbench Bai et al. (2023) and $\infty$-Bench Zhang et al. (2024b). Longbench offers an evaluation on a variety of tasks including question answering, summarization, few-shot learning, mathematical counting, and code completion. The average and $95$% percentile length are 12.8K and 31K tokens, respectively. $\infty$-Bench is designed to test a model’s ability to understand and reason over super long contexts. The average length and the $95$% percentile length are 145.1K and 214K tokens. For more detailed dataset statistics, please refer to Appendix B. We believe that these two benchmarks can comprehensively reflect the capabilities of the model on downstream tasks. For evaluations on the Longbench, we adopt a larger local context size of 3,500 tokens for FocusLLM, consistent with the official setting.
@@
Models. The purpose of our paper is to enable LLMs with limited context length to understand extremely long sequences at a very minimal cost. Therefore, in addition to comparing FocusLLM with the three types of baselines mentioned in Section 3.2, we mainly focused on comparing FocusLLM with recent proposed models capable of processing extremely long streaming inputs. Specifically, StreamingLLM utilizes a sliding window mechanism; InfLLM Xiao et al. (2024) stores processed context into memory units and retrieves it using attention scores; Activation Beacon compresses the preceding text to maintain a smaller context length. CEPE Yen et al. (2024) adopts a small encoder to process long inputs chunk by chunk and feeds the memory to a decoder by cross-attention.
@@
Main Results. The experimental results are displayed in Table 3 and 4. We reference some baseline results from Xiao et al. (2024), which are based on the Vicuna-7B-v1.5 model. Vicuna-7B-v1.5 is based on LLaMA-2-7B but fine-tuned on conversational data. For a fair comparison, we also train a Vicuna version of FocusLLM. For YaRN-128K, we select the version based on Mistral-7B-inst-v0.2, which is stronger than Vicuna. For LongLlama, as they do not have a version based on the Llama2, we directly utilize the officially released model. Details regarding the effective lengths for each method presented in Tables 3 and 4 are elaborated upon in Appendix C.
@@
From the experimental results, we can make the following observations: (1) FocusLLM outperforms all baseline models, achieving the best results on both the relatively shorter benchmark Longbench and the extremely long benchmark $\infty$-Bench. This demonstrates FocusLLM’s capability for effective understanding and reasoning on long sequences and its broad applicability. (2) Different types of baseline models exhibit various shortcomings. For training-free models like PI and NTK, extending the length to 128K comes with a significant sacrifice in performance. Due to the lack of precise understanding of the full context, models that employ sliding window or condensing techniques, such as StreamingLLM and Activation Beacon perform poorly on $\infty$-Bench, with performance nearly approaching zero on some tasks. As for fine-tuned models like LongChat and CEPE, their limitation is the restricted supported length. For example, CEPE struggles to handle lengths beyond 128K effectively Yen et al. (2024). (3) The approaches of length extrapolation and continual training on long inputs, while capable of scaling context length, introduce substantial computational and memory costs. Consequently, this can lead to exceeding GPU memory limits when dealing with long sequences. In contrast, FocusLLM processes the text in chunks and utilizes parallel decoding, which significantly conserves both the memory and time for inference.
In summary, compared to previous context scaling methods, FocusLLM achieves superior performance across a variety of downstream tasks at the same parameter scale, and can be deployed with limited resources.
Table 4: The results of LLaMA2-based models on tasks of LongBench. L_L represents Long Llama and A_B represents Activation Beacon. FocusLLM outperforms recently proposed memory-based and compression-based methods, and maintains attention to all tokens of context.
@@ 4 Further Exploration
In this section, we conduct further explorations of FocusLLM. Initially, we investigate the upper limits of the sequence lengths that FocusLLM can effectively handle. Subsequently, we perform a quantitative analysis of several key parameters within the framework. Ultimately, we undertake ablation studies on the training loss functions to discern their impacts on the model’s overall performance.
@@ 4.1 Scaling to 400K Context
We contend that FocusLLM is capable of processing extremely long sequences. To validate this, we first conduct experiments on the passkey retrieval task Mohtashami and Jaggi (2024). The results, as illustrated in Figure 1, demonstrate that FocusLLM maintains nearly 100% effectiveness at lengths of up to 400K, outperforming all other models. In addition to this, we extended the language modeling experiments introduced in Section 3.2 to 400K, a length at which most models fail to manage effectively. The outcomes of these extended experiments are presented in Figure 6.
To conclude, FocusLLM’s exceptional performance in both passkey retrieval and language modeling tasks, even at the substantial length of 400K tokens [^2] Constrained by hardware, the maximum length we are able to test is 400k tokens., indicates that it can achieve good performance in the vast majority of real-world scenarios, which underscores the effectiveness and versatility of FocusLLM.
Figure 4: Perplextiy on the PG19 dataset of FocusLLM compared to methods PI and NTK. FocusLLM can maintain low perplexity even at a length up to 400K tokens.
Figure 5: Memory usage of FocusLLM compared against previous long context methods. FocusLLM without parallel forwarding exhibits a linear growth pattern.
Figure 6: Comparison of inference time. The time taken by FocusLLM is superior to other long-context methodologies but slightly higher than the standard approach.
@@ 4.2 Memory Footprint and Inference Time
For models that focus on long-form text, aside from training costs, another critical aspect is the memory footprint and inference time, especially as sequence lengths increase. In this section, we compare FocusLLM with several previous long-context methods capable of retaining global information by preserving the cache of all context: Standard (PI/NTK), LongLlama, and CEPE. The results are shown in Figure 6 and Figure 6. As for models like Activation Beacon and StreamingLLM, although they maintain a constant memory footprint by only retaining cache for a fixed window, they struggle with the precise understanding of extremely long texts. Therefore, they are not the primary subjects of comparison.
In the Figure, "FocusLLM with or without parallel" signifies that we process each chunk either concurrently or sequentially. The findings indicate that: (1) Serial processing is good enough in terms of memory usage and inference time. When ample memory resources are available, parallel processing is more efficient. And in practice, the parallel level is also controllable, making our model robust and applicable under most scenarios.(2) Although FocusLLM splits long texts into numerous chunks, resulting in a slightly longer inference time compared to the standard approach, it holds a significant advantage over other long-context methods.
@@ 4.3 Chunk Size
Table 5: Investigations into the training loss function and local context size of FocusLLM (based on LLaMA2-7B). We present the results for representative tasks from LongBench and $\infty$-Bench. For instance, NarrativeQA belongs to Single-Doc QA, while TREC relates to Few-shot learning. The Hyper Params is denoted as (local context size, chunk size).
@@
Repetition Loss only (2K, 2K) 17.05 62.0 12.86 26.64 91.19
@@
4.4 Local Context Size
While the parallel decoding mechanism theoretically allows the model to attend to all tokens, we aim to investigate whether the principle that a larger local context size equates to better performance holds true for FocusLLM as well. The results are shown in the last row of Table 5. As we reduce the local context size from 3.5K to 1K, it can be observed that performance for the majority of tasks experiences a slight decline, with the exception of the passkey retrieval task, which has very clear and concise instructions. This indicates that a larger local context size indeed provides higher quality local information.
@@ 4.5 Ablation Studies
We employ both Continuation Loss and Repetition Loss for the training of FocusLLM. The motivation behind this is to equip the model with the natural language modeling capability while also enhancing its ability to recover global information. Ablation studies on both losses, as detailed in Table 5, reveal that relying solely on the Continuation Loss enables the model to manage some tasks effectively. Nonetheless, for tasks with substantial dependencies on the preceding context, like HotpotQA and Retrieve.PassKey, the model’s efficacy deteriorates. Similarly, while employing the Repetition Loss ensures accurate restatement of the preceding context, the lack of generalizability of generating new tokens leads to a considerable decrease in performance. Therefore, the combined use of both loss functions is crucial for enhancing the performance and generalizability of FocusLLM.
@@ 5.1 Long-context language models
Recent advancements in long-context modeling have seen a surge in innovative approaches that aim to transcend the limitations of transformer architectures. One research direction involves length extrapolation in transformers Peng et al. (2023); Jin et al. (2024), where methods like positional interpolation help models adapt to longer sequences Chen et al. (2023a). However, these techniques often fail to address the distraction issue caused by noisy content within extended texts Tworkowski et al. (2024). Another research branch focus on modifying the attention mechanism or employing compression techniques to maintain long texts within manageable lengths Chevalier et al. (2023); Zhang et al. (2024a). For instance, Xiao et al. (2023) discovered that retaining ‘sink tokens’ in conjunction with a sliding window can achieve smooth streaming output. Zhang et al. (2024a) expanded the context dramatically through compression. However, these methods share a common limitation: they cannot utilize information from all tokens.
@@ 5.2 Memory-enhanced Model
The integration of memory layers within transformer architectures has become a pivotal strategy for enhancing long-context comprehension Bertsch et al. (2024); Tworkowski et al. (2024); Fang et al. (2024). Common methodologies in memory-enhanced models often employ recurrent strategies that iteratively integrate information from the current window into a persistent memory Munkhdalai et al. (2024). Another approach is to initially encode the complete long text into memory tokens, which is then queried in to retrieve pertinent information as needed Xiao et al. (2024). For example, Yen et al. (2024) employ a small encoder to sequentially encode long text segments, followed by the integration of these encoded chunks into a decoder. However, the drawback of such methods is that the memory length does not extrapolate well, and expanding the memory still incurs substantial computational costs. In contrast, FocusLLM offers superior training efficiency and remains effective on exceedingly long texts.
@@ 6 Conclusion
Appendix A More discussion
A.1 Architecture
In this section, we provide an in-depth explanation of the parallel decoding utilized by FocusLLM. As we mentioned in section 2.1.3, the processing of each chunk in FocusLLM can be considered a form of implicit decoding. This is because our goal is not to obtain representations for entire chunks of text, but rather to preserve information about the next token (based on the current chunk) through a special candidate token.
Furthermore, we can observe that the overall structure of FocusLLM shares similarities with the recently proposed concept of the decoder-decoder model architecture Sun et al. (2024), which differs from the traditional encoder-decoder model. Specifically, encoder-decoder models, like T5 Raffel et al. (2020), utilize a bidirectional encoder for input encoding and a unidirectional decoder for output generation. These models necessitate the re-encoding of the entire set of input and output tokens for each subsequent generation step, making them less suitable for autoregressive generation. In contrast, the decoder-decoder model (also FocusLLM), by caching the previously computed key/value vectors, enables the model to reuse them in the current generation step.
A.2 Loss Design
In the loss design of FocusLLM, in addition to using the commonly employed Continuation Loss, we designed Repetition Loss to explicitly train the model to maintain attention to the context. We believe that Repetition Loss is the reason why FocusLLM can achieve good results with a relatively small number of training tokens. Previous work has demonstrated that synthetic data can effectively enhance model capabilities or reduce costs Wu et al. (2024); Li et al. (2022, 2023b, 2024b). This has inspired us to design different types of synthetic data in the future to help FocusLLM quickly learn specific abilities.
Appendix B Details of Benchmarks
B.1 LongBench

### v2

@@ 4 Experiments
In this section, we will conduct a comprehensive evaluation of the effectiveness of FocusLLM, spanning both language modeling and a variety of downstream tasks. We refer readers to Appendix C for detailed experimental settings including hyperparameters due to space constraints.
@@ 4.1 Long-context Language Modeling
In this section, we evaluate FocusLLM on long-context language modeling benchmarks, with text lengths ranging from 4K to 128K tokens.
@@
, and the training-free method StreamingLLM Xiao et al. (2023), which is based on attention sinks. ii) Fine-tuned methods trained on long inputs, such as LongAlpaca-16K Chen et al. (2023b), LongChat-32K Li et al. (2023), and YaRN-128K Peng et al. (2023). iii) Methods with designed structures specifically for long contexts, including AutoCompressor-6K Chevalier et al. (2023), LongLlama Tworkowski et al. (2024) and Activation Beacon Zhang et al. (2024a). For instance, Activation Beacon achieves compression of long texts by training the model to represent the information of a regular text segment with a small number of beacon tokens.
Analysis. The results are presented in Table 1. Here are several observations we can make: (1) Compared to the basic LLaMA-2-7B model and some fine-tuning free methods, our model demonstrates superior performance. When extending the context length from 4K to longer, the perplexity becomes lower, indicating that information from a longer context can be effectively utilized. (2) FocusLLM achieves comparable performance to fine-tuned full-attention methods. This result is notable because our model operates with significantly higher training efficiency. For instance, LongLlama is fine-tuned using 7B tokens with all parameters being trainable. In contrast, FocusLLM uses $1/10$ of the training budget and $1/3$ of the parameters. (3) FocusLLM can maintain language modeling capabilities at lengths much longer than other models while retaining precise comprehension of the entire text. Although models like StreamingLLM and Activation Beacon can still achieve lower perplexity by compressing tokens, they are unable to recover the previous context information, which severely affects their capabilities in downstream tasks. In summary, FocusLLM achieves comparable language modeling performance with a small training cost.
@@
$\infty$-Bench Math.Find 11.71 9.43 13.43 17.14 OOM OOM 6.00 11.14 11.71
@@
Table 2: The results on $\infty$-Bench and LongBench. The models on the right part can process extremely long inputs. On both benchmarks, FocusLLM achieves significant improvements compared to strong baselines.
@@ 4.2 Downstream Tasks
Datasets. To assess the capabilities of FocusLLM in real-world scenarios, we select two widely used datasets: Longbench Bai et al. (2023) and $\infty$-Bench Zhang et al. (2024b). Longbench offers an evaluation on a variety of tasks including question answering, summarization, few-shot learning, mathematical counting, and code completion. $\infty$-Bench is designed to test a model’s ability to understand and reason over super long contexts, with an average length of 145.1K tokens. Thus, the tasks in $\infty$-Bench are well-suited to test whether the model has a precise understanding of long contexts without information loss. For more detailed statistics, please refer to Appendix D. We believe that these two benchmarks can comprehensively reflect the capabilities of the model on downstream tasks.
@@
Models. We select representative models from the three types of baselines mentioned in Section 4.1 for comparison. Additionally, we focus on comparing FocusLLM with recently proposed models capable of processing extremely long streaming inputs. Specifically, StreamingLLM utilizes a sliding window mechanism; InfLLM Xiao et al. (2024) stores processed context into memory units and retrieves it using attention scores; Activation Beacon compresses the preceding text to maintain a smaller context length. CEPE Yen et al. (2024) adopts a small encoder to process long inputs chunk by chunk and feeds the memory to a decoder by cross-attention.
@@
Main Results. The experimental results are displayed in Table 2 and 3. We reference some baseline results from Xiao et al. (2024), which are based on the Vicuna-7B-v1.5 model. Vicuna-7B-v1.5 is based on LLaMA-2-7B but fine-tuned on conversational data. For a fair comparison, we also train a Vicuna version of FocusLLM. For YaRN-128K, we select the version based on Mistral-7B-inst-v0.2, which is stronger than Vicuna. For LongLlama, as they do not have a version based on the Llama2, we directly utilize the officially released model. CEPE and LongLLaMA will experience OOM on $\infty$-Bench due to their substantial memory usage, so we only report their results on LongBench. Since not all models are inherently capable of processing infinite text lengths, we also elaborate the effective lengths for each method presented in Tables 2 and 3 in Appendix E.
@@
From the experimental results, we can make the following comparisons between FocusLLM and previous methods: (1) FocusLLM outperforms all baseline models, achieving the best results on both the relatively shorter benchmark Longbench and the extremely long benchmark $\infty$-Bench. This demonstrates FocusLLM’s capability for effective understanding and reasoning on long sequences and its broad applicability. (2) Different types of baseline models exhibit various shortcomings. For training-free models like PI and NTK, extending the length to 128K comes with a significant sacrifice in performance. Due to the lack of precise understanding of the full context, models that employ sliding window or condensing techniques, such as StreamingLLM and Activation Beacon perform poorly on $\infty$-Bench (see also Appendix F), with performance nearly approaching zero on some tasks. This indicates that they suffer from severe information loss. As for fine-tuned models like LongChat and CEPE, their limitation is the restricted supported length. For example, CEPE struggles to handle lengths beyond 128K effectively Yen et al. (2024). (3) The approaches of length extrapolation and continual training on long inputs, while capable of scaling context, introduce substantial computational and memory costs. In contrast, FocusLLM processes the text in chunks and utilizes parallel decoding, which significantly conserves both the memory and time for inference.
@@
Table 3: The results of LLaMA2-based models on tasks of LongBench. L_L represents Long Llama and A_B represents Activation Beacon. FocusLLM outperforms memory-based and compression-based methods, and maintains attention to all tokens of context.
@@ 5 Further Exploration
5.1 Visualization of Candidate Tokens
To further illustrate how candidate tokens function, we provide a more intuitive explanation by visualizing the information carried by these tokens through attention weight heatmaps when decoding the next token. Due to space limitations, we place the visualization results in Appendix H. We have the following observations: i) In Passkey Retrieval task, the model assigns a high attention weight to one certain candidate token, indicating that this token effectively carry the passkey information from its respective chunk. In contrast, candidate tokens from chunks containing noisy text carry no useful information, resulting in near-zero attention weights. ii) In LongBench NarrativeQA task, the model shows a slightly different pattern, where many candidate tokens receive attention, as multiple chunks’ information may be aggregated for the QA task. The visualization results demonstrate that FocusLLM effectively uses candidate tokens to transmit information from the context while ignoring irrelevant noise.
@@ 5.2 Scaling to 400K Context
We contend that FocusLLM is capable of processing extremely long sequences. To validate this, we first conduct experiments on the passkey retrieval task Mohtashami and Jaggi (2024). The results, as illustrated in Figure 1, demonstrate that FocusLLM maintains nearly 100% effectiveness at lengths of up to 400K [^3] Constrained by hardware, the maximum length we are able to test is 400k tokens., outperforming all other models. We also extended the language modeling experiments introduced in Section 4.1 to 400K, a length at which most models fail to manage effectively. The result is presented in the Appendix G.
Figure 3: FocusLLM exhibits a more efficient growth pattern in memory usage compared to previous methods.
Figure 4: Comparison of inference time. The time taken by FocusLLM is superior to previous methods.
@@ 5.3 Memory Footprint and Inference Time
For models that focus on long texts, aside from training costs, another critical aspect is the memory footprint and inference time. In this section, we compare FocusLLM with several previous long-context methods capable of retaining global information by preserving the cache of all context: Standard (PI/NTK), LongLlama, and CEPE. As for models like Activation Beacon and StreamingLLM, although they maintain a constant memory footprint by only retaining cache for a fixed window, they suffer significant information loss and struggle with the precise understanding of long texts as demonstrated in Section 4.2. Therefore, they are not the primary subjects of comparison.
The results are shown in Figure 5 and Figure 5. FocusLLM with or without parallel indicates whether we process each chunk either concurrently or sequentially. The findings indicate that: (1) When ample memory resources are available, parallel processing is more efficient for FocusLLM. (2) Although FocusLLM splits long texts into numerous chunks, resulting in a slightly longer inference time compared to the standard approach, it still holds a significant advantage over other long-context methods.
@@ 5.4 Chunk Size
Reconstruction Loss only (2K, 2K) 17.05 62.0 12.86 26.64 91.19
@@
Table 4: Investigations into the training loss and local context size of FocusLLM. We present the results for representative tasks from LongBench and $\infty$-Bench. For instance, NarrativeQA belongs to Single-Doc QA, while TREC relates to Few-shot learning. The Hyper Params is denoted as (local context size, chunk size).
@@ 5.5 Ablation Studies
We employ both Continuation Loss and Reconstruction Loss for the training of FocusLLM. The motivation behind this is to equip the model with the natural language modeling capability while also enhancing its ability to recover information. Ablation studies as detailed in Table 4, reveal that relying solely on the Continuation Loss enables the model to manage some tasks effectively. Nonetheless, for tasks with substantial dependencies on the preceding context, like HotpotQA and Retrieve.PassKey, the model’s efficacy deteriorates. Similarly, while employing the Reconstruction Loss ensures accurate restatement of the preceding context, the lack of generalizability of generating new tokens leads to a considerable decrease in performance. Therefore, the combined use of both loss functions is crucial for enhancing the performance and generalizability of FocusLLM.
We also investigate how the local context size influences performance in the last row of Table 4. As we reduce the local context size from 3.5K to 1K, the performance of most tasks experiences a slight decline. This suggests that candidate tokens cannot fully replace the information within the context.
@@ 6.1 Long-context language models
One research direction involves length extrapolation in transformers Peng et al. (2023); Jin et al. (2024), where methods like positional interpolation help models adapt to longer sequences Chen et al. (2023a). However, these techniques often fail to address the distraction issue caused by noisy content within extended texts Tworkowski et al. (2024). Another research branch focus on modifying the attention mechanism or employing compression techniques to maintain long texts within manageable lengths Chevalier et al. (2023); Zhang et al. (2024a). For instance, Xiao et al. (2023) discovered that retaining ‘sink tokens’ in conjunction with a sliding window can achieve smooth streaming output. Zhang et al. (2024a) expanded the context dramatically through compression. However, these methods share a common limitation: they cannot utilize information from all tokens.
@@ 6.2 Memory-enhanced Model
The integration of memory layers within transformer architectures has become a pivotal strategy for enhancing long-context comprehension Bertsch et al. (2024); Tworkowski et al. (2024). Common methodologies in memory-enhanced models often employ recurrent strategies that iteratively integrate information from the current window into a persistent memory Munkhdalai et al. (2024). Another approach is to initially encode the complete long text into memory tokens, which is then queried in to retrieve pertinent information as needed Xiao et al. (2024). For example, Yen et al. (2024) employ a small encoder to sequentially encode long text segments, followed by the integration of these encoded chunks into a decoder. However, the drawback of such methods is that the memory length does not extrapolate well, and expanding the memory still incurs substantial computational costs. In contrast, FocusLLM offers superior training efficiency and remains effective on exceedingly long texts.
@@ 7 Conclusion
8 Limitations
Our research has certain limitations: (1) Due to hardware constraints, our tests were limited to 400K tokens, which does not represent the upper bound of FocusLLM’s capabilities. Future work will explore the full performance potential of FocusLLM and investigate the use of quantization methods to reduce operational costs. (2) While FocusLLM demonstrates exceptional training efficiency, we have observed that training on larger datasets can significantly enhance its generalizability and performance. Therefore, increasing the training data size will be a focus of future research.
Appendix A Efficiency of FocusLLM
Appendix B Details of Training Data
We randomly sampled 80K sequences from RedPajama as our training corpus. Table 5 shows the detailed distribution.
Length 3Ksimilar-$\sim$4K 4Ksimilar-$\sim$6K 6Ksimilar-$\sim$8K Total
Count 30K 16K 34K 80K
Portion 38% 20% 42% 100%
Table 5: Length distribution of training corpus.
Appendix C Experimental Details
We primarily conduct experiments on the LLaMA2-7B-Chat model. The additional trainable parameters mentioned in Section 2 amount to only 2B approximately.
Specifically, we conducted training on a Linux server equipped with 8×A100 GPUs, each with 40GB of memory. The training was carried out for 10,000 steps, equivalent to one epoch of the entire training dataset, using a batch size of 8 and a learning rate of 5e-5 with a linear scheduler. To conserve GPU memory, we employed deepspeed’s zero2_offload optimizing stage. The training process was completed in approximately 20 hours.
For hyper-parameters, during training, the chunk size was randomly selected from the set {64, 128, 256, 1024, 2048}. For the length of tokens injected into each chunk, we set a default of 512 tokens for inference. And we ensured this length did not exceed the chunk size in the training procedure. As a result, the length of injected tokens was $\min\{512,chunk\ size\}$. For evaluations on the Longbench, we adopt a larger local context size of 3,500 tokens for FocusLLM, consistent with the official setting.
Appendix D Details of Benchmarks